One number is 3 times another. If the sum of their reciprocals is 1/5 , find the two numbers.
Answer by stanbon(48558)

One number is 3 times another. If the sum of their reciprocals is 1/5 , find the two numbers.
--------------
Let the numbers be "x" and "3x".
EQUATION
1/x + 1/3x = 1/5
Multiply thru by 15x to get:
15 + 5 = 3x
x = 20/3
3x = 20
==============
Cheers,
Stan H.